OpenStudy (some.random.cool.kid):

What are the domain and range for the cube root function graphed below? Justify your answer in complete sentences.

OpenStudy (some.random.cool.kid):

@Hero

OpenStudy (stefrheart):

Do you have someone else already answering?

OpenStudy (some.random.cool.kid):

No I currently do not

OpenStudy (stefrheart):

Is it okay for me to help? I saw that you commented someone's name

OpenStudy (some.random.cool.kid):

yes hero is a mod here and really brilliant but he doesn't respond to much people So if you think you got this I wont stop you lol

OpenStudy (stefrheart):

Ah I understand :) Do you understand the domain and range?

OpenStudy (some.random.cool.kid):

I know exactly what it is my trouble is finding it on the complex quad graphs so explaining that to me would do me alot of beneficial favors.

OpenStudy (stefrheart):

Well based on this graph there seems to be no restrictions that is stopping the graph in both the x and y axis.

OpenStudy (some.random.cool.kid):

indeed its just asking me to figure the domain and range for it whats bad is that they threw a wild graph at me and I am not sure what any of the points can be... my main struggle is figuring out where it meets I guess ...

OpenStudy (anonymous):

you can take the cubed root of anything

OpenStudy (some.random.cool.kid):

?

OpenStudy (anonymous):

so domain is all real numbers

OpenStudy (stefrheart):

You dont need any points for this graph for finding the domain and range

OpenStudy (stefrheart):

\[(-\infty,\infty)\]

OpenStudy (some.random.cool.kid):

yes ik that but I cant put that for the graph

OpenStudy (some.random.cool.kid):

Im not sure that what its telling me to put when it says find the domain?

OpenStudy (some.random.cool.kid):

thats way to easy ..

OpenStudy (stefrheart):

I know it seems easy but that is the answer... all cube root functions without restrictions have a domain and range of all real numbers. There is nothing else to put besides all real numbers

OpenStudy (some.random.cool.kid):

WOW really thanks! :~> so what about the range?

OpenStudy (some.random.cool.kid):

same i assume?

OpenStudy (stefrheart):

Hahahaha welcome! You are correct, the range is the same! :)

OpenStudy (some.random.cool.kid):

just put the variables for it?

OpenStudy (some.random.cool.kid):

ok

OpenStudy (stefrheart):

Variables?

OpenStudy (some.random.cool.kid):

i mean the -x -y?

OpenStudy (stefrheart):

If you are doing the interval notation is \[(-\infty,\infty)\] and if not then it is xER

OpenStudy (stefrheart):

or yER for range

OpenStudy (some.random.cool.kid):

k thanks bye !D

OpenStudy (stefrheart):

Welcome! Glad to be of help!! :D